# Foreshock Ultra-Low Frequency Waves at Mars: Consequence on the Particle Acceleration Mechanisms at the Martian Bow Shock

### v1

@@ 1 Introduction
Understanding the characteristics of ULF waves in planetary foreshocks is crucial for identifying and comparing physical processes, as well to characterizing particle acceleration mechanisms occurring at planetary bow shocks (e.g., Hoppe & Russell, 1982; Eastwood et al., 2005; Shan et al., 2016, 2018a; Mazelle et al., 2000, 2003; Andrés et al., 2013, 2015; Romanelli et al., 2020; Romanelli & DiBraccio, 2021). Specifically, Hoppe & Russell (1982) investigated the occurrence of foreshock waves at Venus, Earth and Jupiter, using various spacecraft in situ measurements. The authors found an empirical relationship between the magnitude of the interplanetary magnetic field and the observed frequency of these waves. Using Venus Express observations, Shan et al. (2016) conducted a statistical study to identify and characterize ULF waves at the Venus foreshock. Their findings indicated properties of the quasi-monochromatic ULF waves similar to those found in the terrestrial foreshock (see Eastwood et al., 2005). Additionally, Shan et al. (2018a) identified the ULF foreshock wave boundary, suggesting that the backstreaming ions upstream of the Venus bow shock are the primary energy source of these ULF waves. Romanelli et al. (2020) performed a statistical analysis of the main properties of ULF waves at the Hermean foreshock using MErcury Surface, Space ENvironment, GEochemistry, and Ranging (MESSENGER) magnetometer observations. They also found polarization properties similar to those of the quasi-monochromatic “30 s” waves observed at Earth’s foreshock, but with a smaller normalized wave amplitude. Moreover, they estimated the speed of resonant backstreaming protons in the solar wind reference frame.
@@
In contrast with other planets, the task of identifying and characterizing the quasi-monochromatic ULF waves on Mars’s foreshock is significantly more complex due to the concurrent existence of another type of waves known as Proton Cyclotron Waves (PCWs). Due to the lack of an intrinsic global magnetic field, the Martian exosphere is extended beyond the bow shock and is subject to various ionizing mechanisms (see, Chaffin et al., 2014), generating a population of newborn planetary protons (e.g., Yamauchi et al., 2015; Rahmati et al., 2017). The interaction between these newborn protons population and the incoming solar wind population is responsible for the generation of waves observed at a frequency near the local proton cyclotron frequency $f_{ci}$ in the spacecraft reference frame (e.g., Russell et al., 1990; Brain, 2002; Mazelle et al., 2004; Romanelli et al., 2013, 2016; Ruhunusiri et al., 2015, 2016; Liu et al., 2020; Andrés et al., 2020; Romeo et al., 2021). Since these waves have similar properties in the spacecraft frame, such as frequency, polarization and propagation angle with respect to the ambient magnetic field, distinguishing between foreshock ULF waves and PCWs requires a specific methodology. Indeed, to the best of our knowledge, there is only one reported event of foreshock ULF wave activity at Mars. Using magnetic field and plasma observations from the Mars Atmosphere and Volatile EvolutioN (MAVEN) mission, Shan et al. (2020) reported an event of ULF waves when MAVEN was connected to the Martian bow shock. Spectral analysis revealed that the frequency of this wave event was approximately 0.040 Hz in the spacecraft frame, approximately a factor two lower than the local proton cyclotron frequency $f_{ci}\sim$ 0.088 Hz. Additionally, the authors reported a propagation angle of about 34 relative to the mean magnetic field, with a left-hand elliptical polarization in the spacecraft frame, and a phase speed much lower than the solar wind velocity. Numerically, Jarvinen et al. (2022) investigated the solar wind interaction with Mars using a global three-dimensional hybrid model. The authors found a well-developed, extensive ion foreshock and the presence of large-scale ULF waves in two specific regions. In the so-called near region, the wave period is 71–83 s and the propagation is nearly perpendicular to the magnetic field with both left- and right-handed wave polarization. On the other hand, a far region where the wave period is 25–28 s, the propagation angle varies between 20 and 50, and the waves present a left-handed polarization with respect to the ambient magnetic field.
@@ 2 MAVEN MAG Observations and Detection Criteria
The MAVEN mission was launched on November 18, 2013, and entered Mars’ orbit on September 21, 2014. Its primary objectives are to investigate Mars’ upper atmosphere, ionosphere, and their interactions with the solar wind. MAVEN’s orbit features a nominal periapsis altitude of 150 km, an apoapsis altitude of 6220 km, and an orbital period of approximately 4.5 h (Jakosky et al., 2015). The significant apoapsis altitude enable the investigation of plasma properties upstream of Mars bow shock, particularly in the Martian foreshock region. In the present work, we have analyzed MAVEN MAG observations with 32 Hz sampling cadence (Connerney et al., 2015) from November 17, 2014 until November 30, 2016 upstream from the Martian bow shock. We analyzed time intervals when MAVEN was magnetically connected to the Martian bow shock. To achieve this, we separated the time series into 255-s events, since this specific time duration for each event was enough to identify the foreshock ULF waves (see, Hoppe & Russell, 1982; Shan et al., 2020). For each event, we calculated the foreshock coordinates (Greenstadt & Baum, 1986) using the spacecraft’s position, the local mean magnetic field, and a Martian bow shock model (Gruesbeck et al., 2018), determining whether MAVEN was within the foreshock region. It is worth mentioning, that this technique have been extensively used in the literature (e.g., Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Kajdič et al., 2017; Romanelli et al., 2020; Romanelli & DiBraccio, 2021).
@@
    We defined an event with ULF wave activity if the PSD showed an increase within a frequency interval centered around a specific frequency $f_{0}$ compared to two adjacent windows each with a width 0.2 $f_{0}$. Given the potential variability in the wave frequency ($f_{0}$) (see, Romanelli et al., 2020), we applied the following condition for each frequency between 0.007 and 0.090 Hz,
@@
    Finally, we identified for the frequency $f_{0}$ where the above condition was maximized, i.e., where the ratio of integrals as a function of the frequency $f_{0}$ was maximum, and we identified this frequency as the ULF foreshock wave frequency $f_{fw}$.
It is worth mentioning that the frequency range used for $f_{0}$ is based on the magnetic field magnitude and the expected range of frequencies for these waves at Mars’ heliocentric distance (see Figure 1 in, Hoppe & Russell, 1982). Also, as we discussed in the Introduction, it is important to note that the foreshock ULF wave frequency range may overlaps with the PCWs frequency range in the Martian foreshock. We shall back to this point in Section 4.2.
@@ 3 Foreshock ULF waves: a case study on January 20, 2015
Figure 1 presents a foreshock ULF wave event identified using the detection criteria described in Section 2. Figure 1 (a-d) show the magnetic field components and magnitude as a function of time for a 255-s event seen by MAVEN on DOY 20 2015 at 01:39:46.85 UT. The magnetic field components are shown in the Mars-centered Solar Orbital (MSO) coordinates, where the ^$\hat{\bf}x$ direction points toward the Sun, the ^$\hat{\bf}z$ direction is perpendicular to Mars’s orbital plane (positive toward the ecliptic north) and the ^$\hat{\bf}y$ axis completes the right-handed system. The three magnetic field components exhibit oscillations with a well-defined frequency. Figure 1 (e) shows the PSD for the total magnetic field as a function of the frequency. Our method identified a clear peak at 0.0541 Hz, corresponds to the foreshock wave frequency $f_{fw}$. It is noteworthy that the PCW frequency $f_{ci}$ for this event is 0.0758 Hz, determined by the magnetic field magnitude (approximately $\sim$ 5 nT) since $f_{ci}=q|{\bf B}|/m$. Additionally, a slight increase in power at $f\sim 1$ Hz, indicating the presence of whistler waves upstream of the planetary bow shock, commonly referred to as “1 Hz” waves (see, e.g., Jin et al., 2024). These waves are typically found in planetary foreshocks, suggesting the employed selection criteria are appropriate.
@@
Figure 3: (a) Foreshock ULF wave frequencies $f_{fw}$ as a function of PCW frequencies $f_{ci}$ along with the distribution of values. The dashed-blue line corresponds to the best linear fit and the dotted-gray line corresponds to $f_{fw}=f_{ci}$ and the shade gray area is limited by the lines $f_{fw}=1.15f_{ci}$ and $f_{fw}=0.85f_{ci}$, respectively. (b) Figure in the same format than (a), but removing the potential PCW activity.
@@ 4.1 Minimum Variance Analysis Results
4.2 Foreshock ULF wave vs PCW wave frequencies
Figure 3 (a) shows ULF foreshock wave frequencies $f_{fw}$ as a function of PCW frequencies $f_{ci}$, along with their distribution. The dotted gray line represents $f_{fw}=f_{ci}$ and the shaded gray area is bounded by the lines $f_{fw}=1.15f_{ci}$ and $f_{fw}=0.85f_{ci}$. The dashed blue line indicates the best linear fit, given by $f_{fw}=\alpha f_{ci}$ with $\alpha=0.67\pm 0.05$. As we discussed in the Introduction, in contrast with other planets, as Mercury (e.g., Romanelli & DiBraccio, 2021) or the Earth (e.g., Andrés et al., 2015), the Martian foreshock exhibits the coexistence of at least two types of waves. One type includes the ULF foreshock waves analyzed in the present work, while the other consists of proton cyclotron waves, which are observed at frequencies near the local proton cyclotron frequency $f_{ci}$. Since, these PCWs present similar wave properties than the ULF foreshock waves, it is not possible to separate them using the single spacecraft techniques employed in this study. Specifically, within the region bounded by $f_{fw}=1.15f_{ci}$ and $f_{fw}0.85f_{ci}$, it is unclear whether we are observing ULF foreshock waves or PCWs. Therefore, in Figure 3 (b) we present the ULF foreshock wave frequencies as a function of the PCW wave frequencies in the same format as panel (a), but excluding potential PCW activity, i.e., those events with frequencies around the proton cyclotron frequency. After removing the potential PCWs, the best linear fit is $f_{fw}=\alpha f_{ci}$, with $\alpha=0.56\pm 0.05$.
@@ 4.3 Particle acceleration mechanisms
The foreshock ULF wave properties discussed in the previous Sections can be used to estimate properties of the backstreaming ions responsible for generating these waves (Gary, 1993; Meziane & d’Uston, 1998; Mazelle et al., 2003; Meziane et al., 2007). Specifically, two commonly used parameters can be estimated, the ratio of the particle velocity parallel to the mean magnetic field to the solar wind velocity ($V_{r}/V_{sw}$) and the ratio of the guiding center velocity of a backstreaming particle to the solar wind velocity ($V_{gc}/V_{sw}$). For a detailed derivation of these parameter, readers are referred to Meziane & d’Uston (1998) and Romanelli et al. (2020) and the reference therein. Briefly, using the cyclotron resonance condition between a backstreaming particle and a linear right-hand wave (in the plasma reference frame) and the Doppler shift equation between the plasma and the spacecraft reference frames, both parameters can be expressed as,
@@
where $\theta_{Bx}$ is the angle between the ambient magnetic field and the mean solar wind velocity, and $a=0.15$ (see, Gary, 1978). It is worth mentioning that the statistical results presented here do not show a strong dependence on the value of $a$ for the potential values that $a$ can take (see, Romanelli et al., 2020; Mazelle et al., 2003).
Figure 4 (a-b) present the distributions of $V_{r}/V_{sw}$ and $V_{gc}/V_{sw}$, respectively. The shaded distributions correspond to the events present in Figure 3 (a) and the bold distributions correspond to the event in Figure 3 (b). Our statistical analysis for all the cases shows that means values are $V_{r}/V_{sw}=1.14\pm 0.38$ and $V_{gc}/V_{sw}=2.02\pm 0.4$. Notably, after removing the potential PCW activity, we obtain similar mean results with $V_{r}/V_{sw}=1.33\pm 0.40$ and $V_{gc}/V_{sw}=2.2\pm 0.43$. Moreover, these observational results fall within the range of previously reported ratios for different planetary foreshocks (see, Hoppe & Russell, 1982; Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Shan et al., 2018b; Romanelli et al., 2020; Romanelli & DiBraccio, 2021).
@@ 5 Discussion and Conclusions
To the best of our knowledge, our study presents the first statistical analysis of ultra-low frequency waves at the Martian foreshock. Utilizing over two years of data from the Magnetometer instrument onboard the MAVEN spacecraft, we focused on periods when MAVEN was magnetically connected to Mars’ bow shock. By applying a specific criterion for wave identification, we determined that the foreshock ULF waves exhibit moderate wave amplitudes and frequencies typically ranging from approximately 0.008 Hz to 0.086 Hz. Based on the MVA, these waves predominantly display circular or elliptical left-handed polarization in the spacecraft reference frame, consistent with a distribution of $\lambda_{1}/\lambda_{2}$ ratios between 1 and approximately 3.
@@
As discussed in the Introduction, the Martian foreshock contains ULF waves and another type of wave known as proton cyclotron waves, which share similar frequency range and polarization properties. Unfortunately, with the techniques and observations used in this study, we cannot completely exclude the presence of PCW activity in our analysis. However, it is important to note that excluding events with frequencies within the expected range for PCWs (i.e., foreshock frequencies between $f_{fw}=0.85f_{ci}$ and $f_{fw}=1.15f_{ci}$), does not significantly affect the statistical properties of the waves or their implications for the acceleration models. An interesting result is that removing potential PCWs events changes the slope $\alpha$ between the foreshock ULF wave frequency $f_{fw}$ and the PCW frequency $f_{ci}$ changes from $(0.67\pm 0.05)$ to $(0.56\pm 0.05)$. Romanelli et al. (2020) conducted a statistical analysis of ULF wave activity in Mercury’s foreshock using high-resolution magnetic field measurements from MESSENGER. The authors found properties similar to those analyzed here and identified a clear relationship between the observed frequency $f_{fw}$ and the local magnetic field magnitude $B$. Specifically, they determined that $f_{fw}(\text{Hz})=(0.00796\pm 0.00170)~{}|B|~{}(\text{nT})$, which corresponds to $f_{fw}=(0.52\pm 0.09)f_{ci}$. Notably, their findings agree within uncertainties with earlier observational results reported by Hoppe & Russell (1982). Using data from the foreshocks of Jupiter, Earth, Venus and Mercury, Hoppe & Russell (1982) established an empirical relationship between the observed ULF wave frequency and the local magnetic field strength $f_{fw}(\text{Hz})=0.0058~{}|B|~{}(\text{nT})$. Our relationship between the observed wave frequency and the local magnetic field strongly agrees with these previous statistical results from other planetary foreshock, particularly after removing the potential PCW contamination.
@@
Using the foreshock ULF wave properties obtained from the minimum variance analysis, we estimated the characteristics of the backstreaming ions responsible for generating the ULF waves (e.g., see Gary, 1993). In particular, we estimated the ratio between the particle velocity parallel to the mean magnetic field and the solar wind velocity $V_{r}/V_{sw}$, and the ratio between the guiding center velocity of a backstreaming particle and the solar wind velocity $V_{gc}/V_{sw}$. In the past, these two parameters have been used to limit the potential mechanism present in planetary bow shocks (see, Hoppe & Russell, 1982; Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Shan et al., 2018b; Romanelli et al., 2020; Romanelli & DiBraccio, 2021). For the Earth’s foreshock, Meziane & d’Uston (1998) and Andrés et al. (2015) reported $\langle V_{gc}/V_{sw}\rangle=1.05\pm 0.01(\theta_{Bx}=45^{\circ})$, $\langle V_{gc}/V_{sw}\rangle=1.11\pm 0.04$ and $\langle V_{gc}/V_{sw}\rangle=1.68$ (for the field-aligned beam boundary), respectively. More recently, Romanelli et al. (2020) reported $\langle V_{gc}/V_{sw}\rangle=0.76\pm 0.27$ for Mercury, and Shan et al. (2018b) reported $\langle V_{gc}/V_{sw}\rangle=1.07~{}(\theta_{Bx}=36^{\circ})$ for Venus. Our results are compatible with these previous observational findings. However, we are obtaining slightly larger values of $\langle V_{gc}/V_{sw}\rangle$ due to the probable inclusion of mix contributions from different cone angle. Additionally, as was pointed out by Romanelli et al. (2020), we cannot ruled out a potential dependence on the size of the bow shock, the tangential convective electric field, or even the heliocentric distance. For the $V_{r}/V_{sw}$ ratio, previously observational results are reported as 1.2-2.2 for Mercury, 1.7/1.9 for Venus, 2.5$\pm$0.3 for Earth and 2.1/2.3 for Jupiter (see, Hoppe & Russell, 1982; Romanelli et al., 2020). Our distribution of $V_{r}/V_{sw}$ encompasses these previous observational results and is consistent with resonant conditions for the backstreaming distributions with velocities varying between approximately 0.5 and 2 times the solar wind velocity. The similarity of this range to previous planetary foreshock observations may indicates the presence of an universal acceleration process acting in planetary shocks and responsible for this particular energetic backstreaming population. Finally, it is worth mentioning that a deep and conclusive analysis on the possible acceleration mechanisms responsible for the backstreaming distributions necessarily involves the use of particle observations in the Martian foreshock and is beyond the scope of the present work.
@@
Our results reveal that Martian ULF foreshock waves share similarities with those observed at other planets, showing a potential universality on their characteristics. The observed linear relationship between the frequency and the magnetic field strength further corroborates the idea that these waves are produced by resonant instabilities, despite the presence of newborn planetary ions and PCWs. Our assessment of backstreaming proton energy, aligning with values found in other planetary foreshocks, indicates that gyro-resonance instabilities significantly contribute to wave generation. This raises the intriguing question of whether one type of plasma instability might dominate over another within the turbulent plasma environment (Ruhunusiri et al., 2017; Halekas et al., 2020; Andrés et al., 2020, 2021; Romanelli et al., 2022, 2024). Mars, with its unique heliospheric position and extended hydrogen exosphere, provides an exceptional platform for investigating these phenomena, offering insights into additional plasma instabilities that could overshadow traditional planetary foreshock instabilities. These findings emphasize the significance of Mars as a natural laboratory for advancing our understanding of planetary foreshock dynamics and their broader implications for space plasma physics.

### v2

@@ 1 Introduction
Understanding the characteristics of ULF waves in planetary foreshocks is crucial for identifying and comparing physical processes, as well as characterizing the particle acceleration mechanisms occurring at planetary bow shocks (e.g., Hoppe & Russell, 1982; Eastwood et al., 2005; Shan et al., 2016, 2018a; Mazelle et al., 2000, 2003; Andrés et al., 2013, 2015; Romanelli et al., 2020; Romanelli & DiBraccio, 2021). Specifically, Hoppe & Russell (1982) investigated the occurrence of foreshock waves at Venus, Earth and Jupiter, using various spacecraft in situ measurements. The authors found an empirical relationship between the magnitude of the interplanetary magnetic field and the observed frequency of these waves. Using Venus Express observations, Shan et al. (2016) conducted a statistical study to identify and characterize ULF waves at the Venus foreshock. Their findings indicated properties of the quasi-monochromatic ULF waves similar to those found in the terrestrial foreshock (see Eastwood et al., 2005). Additionally, Shan et al. (2018a) identified the ULF foreshock wave boundary, suggesting that the backstreaming ions upstream of the Venus bow shock are the primary energy source of these ULF waves. Romanelli et al. (2020) performed a statistical analysis of the main properties of ULF waves at the Hermean foreshock using MErcury Surface, Space ENvironment, GEochemistry, and Ranging (MESSENGER) magnetometer observations. They also found polarization properties similar to those of the quasi-monochromatic “30 s” waves observed at Earth’s foreshock, but with a smaller normalized wave amplitude. Moreover, they estimated the speed of resonant backstreaming protons in the solar wind reference frame.
@@
In contrast with other planets, the task of identifying and characterizing the quasi-monochromatic ULF waves on Mars’s foreshock is significantly more complex due to the concurrent existence of another type of waves known as Proton Cyclotron Waves (PCWs). Due to the lack of an intrinsic global magnetic field, the Martian exosphere is extended beyond the bow shock and is subject to various ionizing mechanisms (see, Chaffin et al., 2014), generating a population of newborn planetary protons (e.g., Yamauchi et al., 2015; Rahmati et al., 2017). The interaction between these newborn protons population and the incoming solar wind population is responsible for the generation of waves observed at a frequency near the local proton cyclotron frequency $f_{ci}$ in the spacecraft reference frame (e.g., Russell et al., 1990; Brain, 2002; Mazelle et al., 2004; Romanelli et al., 2013, 2016; Ruhunusiri et al., 2015, 2016; Liu et al., 2020; Andrés et al., 2020; Romeo et al., 2021). Since these waves have similar properties in the spacecraft frame, such as frequency, polarization, and propagation angle with respect to the ambient magnetic field, distinguishing between foreshock ULF waves and PCWs requires a specific methodology. Indeed, to the best of our knowledge, there has been only one reported event of foreshock ULF wave activity at Mars. Using magnetic field and plasma observations from the Mars Atmosphere and Volatile EvolutioN (MAVEN) mission, Shan et al. (2020) reported an event of ULF waves when MAVEN was connected to the Martian bow shock. Spectral analysis revealed that the frequency of this wave event was approximately 0.040 Hz in the spacecraft frame, approximately a factor two lower than the local proton cyclotron frequency $f_{ci}\sim$ 0.088 Hz. Additionally, the authors reported a propagation angle of about 34 relative to the mean magnetic field, with a left-hand elliptical polarization in the spacecraft frame, and a phase speed much lower than the solar wind velocity. Numerically, Jarvinen et al. (2022) investigated the solar wind interaction with Mars using a global three-dimensional hybrid model. The authors found a well-developed, extensive ion foreshock and the presence of large-scale ULF waves in two specific regions. In the so-called near region, the wave period is 71–83 s and the propagation is nearly perpendicular to the magnetic field with both left- and right-handed wave polarization. On the other hand, a far region where the wave period is 25–28 s, the propagation angle varies between 20 and 50, and the waves present a left-handed polarization with respect to the ambient magnetic field.
@@ 2 MAVEN MAG Observations and Detection Criteria
The MAVEN mission was launched on November 18, 2013, and entered Mars’ orbit on September 21, 2014. Its primary objectives are to investigate Mars’ upper atmosphere, ionosphere, and their interactions with the solar wind. MAVEN’s orbit features a nominal periapsis altitude of 150 km, an apoapsis altitude of 6220 km, and an orbital period of approximately 4.5 h (Jakosky et al., 2015). The significant apoapsis altitude enable the investigation of plasma properties upstream of Mars bow shock, particularly in the Martian foreshock region. In the present work, we have analyzed MAVEN MAG observations with 32 Hz sampling cadence (Connerney et al., 2015) from November 17, 2014 until November 30, 2016 upstream from the Martian bow shock. To identify ULF waves in Mars’ foreshock region, first we analyzed time intervals when MAVEN was magnetically connected to the Martian bow shock. To achieve this, we separated the time series into 255-s events and calculated the foreshock coordinates (Greenstadt & Baum, 1986) using the spacecraft’s position, the local mean magnetic field, and a Martian bow shock model (Gruesbeck et al., 2018). This technique has been extensively used in the literature to estimate the spacecraft location with respect to a planetary foreshock (e.g., Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Kajdič et al., 2017; Romanelli et al., 2020; Romanelli & DiBraccio, 2021). The specific time duration for each events was selected to include the potential frequencies where we expect foreshock ULF waves (see, Hoppe & Russell, 1982; Shan et al., 2020). More specifically, we looked for foreshock waves based on the magnetic field magnitude and the expected range of frequencies for these waves at Mars’ heliocentric distance (see Figure 1 in, Hoppe & Russell, 1982).
@@
    We defined an event with ULF wave activity if the PSD showed an increase within a frequency interval centered around a specific frequency $f_{0}$ compared to two adjacent windows each with a width $\Delta f=0.2f_{0}$. Given the potential variability in the wave frequency ($f_{0}$) (see, Romanelli et al., 2020), we applied the following condition for each frequency between 0.007 and 0.090 Hz,
@@
    Finally, we identified the frequency $f_{0}$ where the above condition was maximized, i.e., where the ratio of integrals as a function of the frequency $f_{0}$ was maximum, and we denote this frequency as $f_{id}$.
As we discussed in the Introduction, it is important to note that the identified wave frequency $f_{id}$ may falls within the PCWs frequency range in the Martian foreshock. We shall back to this point in Section 4.1. Also, it is worth mentioning that we used a width of contiguous windows $\Delta f=0.2f_{0}$, a typical value for waves identification (Romanelli et al., 2016; Andrés et al., 2020; Romeo et al., 2021). However, when we range $\Delta f$ between $0.1f_{0}$ and $0.3f$, we observed a negligible impact on our statistical results.
@@ 3 Foreshock ULF waves: a case study on January 20, 2015
Figure 1 presents a foreshock ULF wave event identified using the detection criteria described in Section 2. Figure 1 (a-d) show the magnetic field components and magnitude as a function of time for a 255-s event seen by MAVEN on DOY 20 2015 at 01:39:46.85 UT. The magnetic field components are shown in the Mars-centered Solar Orbital (MSO) coordinates, where the $\hat{\bf x}$ direction points toward the Sun, the $\hat{\bf z}$ direction is perpendicular to Mars’s orbital plane (positive toward the ecliptic north) and the $\hat{\bf y}$ axis completes the right-handed system. The three magnetic field components exhibit oscillations with a well-defined frequency. Figure 1 (e) shows the PSD for the total magnetic field as a function of the frequency. Our method identified a clear peak at 0.0541 Hz, corresponds to the foreshock wave frequency $f_{fw}$. It is noteworthy that the PCW frequency $f_{ci}$ for this event is 0.0758 Hz, determined by the magnetic field magnitude (approximately $\sim$ 5 nT) since $f_{ci}=q|{\bf B}|/m$. Additionally, a slight increase in power at $f\sim 1$ Hz, indicating the presence of whistler waves upstream of the planetary bow shock, commonly referred to as “1 Hz” waves (see, e.g., Jin et al., 2024). These waves are typically found in planetary foreshocks, suggesting the employed selection criteria are appropriate.
@@
Figure 2: Identified frequencies $f_{id}$ as a function of PCW frequencies $f_{ci}$. The colorbar corresponds to the number of events per bin.
Figure 3: Foreshock ULF wave frequencies $f_{fw}$ as a function of PCW frequencies $f_{ci}$. The dot-gray line corresponds to $f_{fw}=f_{ci}$. The dashed-blue line corresponds to the best linear fit for the foreshock ULF wave events.
@@ 4 Statistical Results
4.1 Foreshock ULF wave frequencies
Figure 2 shows the frequencies identified $f_{id}$ as a function of PCW frequencies $f_{ci}$, for all events. The dotted gray line represents $f_{id}=f_{ci}$. As we discussed in the Introduction, in contrast with other planets, such as Mercury (e.g., Romanelli & DiBraccio, 2021) or Earth (e.g., Andrés et al., 2015), the Martian foreshock exhibits the coexistence of at least two types of waves. One type includes the ULF foreshock waves analyzed in the present work, while the other consists of proton cyclotron waves, which are observed at frequencies close to the local proton cyclotron frequency $f_{ci}$. Since these PCWs present wave properties similar to those of the ULF foreshock waves (see, Hoppe & Russell, 1982; Romanelli & DiBraccio, 2021), it is challenging to identifying them using single spacecraft techniques. Specifically, within the region bounded by $f_{id}\leq 0.85f_{ci}$ (see Delva et al., 2008), it is unclear whether we observe PCWs in the foreshock region or ULF foreshock waves. Therefore, to ensure that we are studying only foreshock ULF waves, we excluded those events from our dataset.
[…]
@@ 4.3 Particle acceleration mechanisms
The foreshock ULF wave properties discussed in the previous Sections can be used to estimate properties of the backstreaming ions responsible for generating these waves (Gary, 1993; Meziane & d’Uston, 1998; Mazelle et al., 2003; Meziane et al., 2007). Specifically, two commonly used parameters can be estimated, the ratio of the particle velocity parallel to the mean magnetic field to the solar wind velocity ($V_{r}/V_{sw}$) and the ratio of the guiding center velocity of a backstreaming particle to the solar wind velocity ($V_{gc}/V_{sw}$). For a detailed derivation of these parameters, the reader is referred to Meziane & d’Uston (1998) and Romanelli et al. (2020) and the references therein. Briefly, using the cyclotron resonance condition between a backstreaming particle and a linear right-hand wave (in the plasma reference frame) and the Doppler shift equation between the plasma and the spacecraft reference frames, both parameters can be expressed as,
@@
where $\theta_{Bx}$ is the angle between the ambient magnetic field and the mean solar wind velocity, and $a\equiv\omega/\Omega_{p}$, where $\omega$ is the wave frequency in the solar wind rest frame and $\Omega_{p}$ is the proton gyro-frequency. While the ${V_{r}}/{V_{sw}}$ ratio is derived in the plasma reference frame, the ${V_{r}}/{V_{sw}}$ ratio is expressed in the spacecraft reference frame.
Figure 5 (a-b) shows the distributions of $V_{r}/V_{sw}$ and $V_{gc}/V_{sw}$, respectively. Our statistical analysis shows that the mean values are $V_{r}/V_{sw}=1.33\pm 0.40$ and $V_{gc}/V_{sw}=2.20\pm 0.43$. It is worth mentioning that we used a value close to what is observed in the terrestrial foreshock $a=0.15$. However, it should be noted that the mean, median, and standard deviation of $V_{r}/V_{sw}$ and $V_{gc}/V_{sw}$ do not show strong variability when we range $a$ between 0.05 and 0.25 (see also, Romanelli et al., 2020). Our observational results fall within the range of previously reported ratios for different planetary foreshocks (see, Hoppe & Russell, 1982; Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Shan et al., 2018b; Romanelli et al., 2020; Romanelli & DiBraccio, 2021).
@@ 5 Discussion and Conclusions
To the best of our knowledge, our study presents the first statistical analysis of ultra-low frequency waves at the Martian foreshock. Using over two years of data from the Magnetometer instrument onboard the MAVEN spacecraft, we focused on periods when MAVEN was magnetically connected to Mars’ bow shock. By applying a specific criterion for wave identification, we determined that the foreshock ULF waves exhibit moderate wave amplitudes and frequencies typically ranging from approximately 0.008 Hz to 0.086 Hz. Based on the MVA, these waves predominantly display circular or elliptical left-handed polarization in the spacecraft reference frame, consistent with a distribution of $\lambda_{1}/\lambda_{2}$ ratios between 1 and approximately 3.
@@
As discussed in the Introduction, the Martian foreshock contains ULF waves and another type of wave known as proton cyclotron waves, which share similar frequency range and polarization properties. Unfortunately, with the techniques and observations used in this study, we cannot completely exclude the presence of PCW activity in our analysis. However, it is important to note that excluding events with frequencies within the expected range for PCWs (i.e., foreshock frequencies between $f_{fw}=0.85f_{ci}$ and $f_{fw}=1.15f_{ci}$), does not significantly affect the statistical properties of the waves or their implications for the acceleration models. After removing potential PCWs events, the slope $\alpha$ between the foreshock ULF wave frequency $f_{fw}$ and the PCW frequency $f_{ci}$ is $(0.7\pm 0.15)$. Romanelli et al. (2020) conducted a statistical analysis of ULF wave activity in Mercury’s foreshock using high-resolution magnetic field measurements from MESSENGER. The authors found properties similar to those analyzed here and identified a clear relationship between the observed frequency $f_{fw}$ and the local magnetic field magnitude $B$. Specifically, they determined that $f_{fw}(\text{Hz})=(0.00796\pm 0.00170)~{}|B|~{}(\text{nT})$, which corresponds to $f_{fw}=(0.52\pm 0.11)f_{ci}$. Notably, their findings agree within uncertainties with earlier observational results reported by Hoppe & Russell (1982).
@@
Using the foreshock ULF wave properties obtained from the minimum variance analysis, we estimated the characteristics of the backstreaming ions responsible for generating the ULF waves (e.g., see Gary, 1993). In particular, we estimated the ratio between the particle velocity parallel to the mean magnetic field and the solar wind velocity $V_{r}/V_{sw}$, and the ratio between the guiding center velocity of a backstreaming particle and the solar wind velocity $V_{gc}/V_{sw}$. In the past, these two parameters have been used to limit the potential mechanism present in planetary bow shocks (see, Hoppe & Russell, 1982; Meziane & d’Uston, 1998; Andrés et al., 2013, 2015; Shan et al., 2018b; Romanelli et al., 2020; Romanelli & DiBraccio, 2021). For the Earth’s foreshock, Meziane & d’Uston (1998) and Andrés et al. (2015) reported $\langle V_{gc}/V_{sw}\rangle=1.05\pm 0.01(\theta_{Bx}=45^{\circ})$, $\langle V_{gc}/V_{sw}\rangle=1.11\pm 0.04$ and $\langle V_{gc}/V_{sw}\rangle=1.68$ (for the field-aligned beam boundary), respectively. More recently, Romanelli et al. (2020) reported $\langle V_{gc}/V_{sw}\rangle=0.76\pm 0.27$ for Mercury, and Shan et al. (2018b) reported $\langle V_{gc}/V_{sw}\rangle=1.07~{}(\theta_{Bx}=36^{\circ})$ for Venus. Our results are compatible with these previous observational findings. However, we are obtaining slightly larger values of $\langle V_{gc}/V_{sw}\rangle$ due to the probable inclusion of mix contributions from different cone angle. For the $V_{r}/V_{sw}$ ratio, previously observational results are reported as 1.2-2.2 for Mercury, 1.7/1.9 for Venus, 2.5$\pm$0.3 for Earth and 2.1/2.3 for Jupiter (see, Hoppe & Russell, 1982; Romanelli et al., 2020). Our distribution of $V_{r}/V_{sw}$ encompasses these previous observational results and is consistent with resonant conditions for the backstreaming distributions with velocities varying between approximately 0.5 and 2 times the solar wind velocity. The similarity of this range to previous planetary foreshock observations may indicates the presence of an universal acceleration process acting in planetary shocks and responsible for this particular energetic backstreaming population. Finally, it is worth mentioning that a deep and conclusive analysis on the possible acceleration mechanisms responsible for the backstreaming distributions necessarily involves the use of particle observations in the Martian foreshock and is beyond the scope of the present work.
@@
Our results reveal that Martian ULF foreshock waves share similarities with those observed at other planets, showing a potential universality on their characteristics. The observed linear relationship between the frequency and the magnetic field strength further corroborates the idea that these waves are produced by resonant instabilities, despite the presence of newborn planetary ions and PCWs. Our assessment of backstreaming proton energy, aligning with values found in other planetary foreshocks, indicates that gyro-resonance instabilities significantly contribute to wave generation. This raises the intriguing question of whether one type of plasma instability might dominate over another within the turbulent plasma environment (Ruhunusiri et al., 2017; Halekas et al., 2020; Andrés et al., 2020, 2021; Romanelli et al., 2022, 2024, 2024). Mars, with its unique heliospheric position and extended hydrogen exosphere, provides an exceptional platform for investigating these phenomena, offering insights into additional plasma instabilities that could overshadow traditional planetary foreshock instabilities. These findings emphasize the significance of Mars as a natural laboratory for advancing our understanding of planetary foreshock dynamics and their broader implications for space plasma physics.